# citations.bib

### v1

@@ D.2 Ratio of mean squared errors
Given an $r\times q$ matrix $\bm{U}$ with orthonormal columns, and an $r$-dimensional vector $\bm{b}$, suppose that we want to estimate $\bm{U}^{\mathsf{T}}\bm{b}$ using matrix multiplication estimator (see Theorem A.1) and $\operatornamewithlimits{argmin}_{\bm{y}}\left\lVert\bm{U}\bm{y}-\bm{b}\right\rVert$ using a regression estimator (see Theorem A.4). Given a sampling distribution $\mathcal{P}$ over $[r]$ and a fixed number of samples $m$, denote $\bm{X}_{m}^{\textup{M}}(\mathcal{P})$, $\bm{X}_{m}^{\textup{R}}(\mathcal{P})$ to be the matrix multiplication estimator and regression estimator for $\bm{U}^{\mathsf{T}}\bm{b}$, respectively.
Given two sampling distributions $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$, denote $\gamma_{1}$ and $\gamma_{2}$ to be the values of $\gamma$ as defined in Eq. 2.3 with respect to distributions $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$, respectively. Fix the number of samples $m\in\mathbb{N}$. Then, we have the following results.
@@
Let $\bm{S}$ be an $m\times r$ sketch matrix (for sampling with replacement) as defined in Eq. A.2 with respect to the distribution $\mathcal{P}$. Then, the matrix multiplication estimator is $\bm{X}_{m}^{\textup{M}}(\mathcal{P})=\bm{U}^{\mathsf{T}}\bm{S}^{\mathsf{T}}\bm{S}\bm{b}$, while the regression estimator is $\bm{X}_{m}^{\textup{R}}(\mathcal{P})=\operatornamewithlimits{argmin}_{\bm{y}}\lVert\bm{S}\bm{U}\bm{y}-\bm{S}\bm{b}\rVert$.
1. Matrix multiplication: Since $\mathbb{E}[\|\bm{X}_{m}^{\textup{M}}(\mathcal{P})-\bm{U}^{\mathsf{T}}\bm{b}\|^{2}]$ is the variance of $\bm{X}_{m}^{\textup{M}}$ using $m$ samples, from Eq. A.4, we have
@@
for $i=1,2$, from which we obtain Eq. D.7.
2. Regression: Observe that $\operatornamewithlimits{argmin}_{\bm{y}}\lVert\bm{U}\bm{y}-\bm{b}\rVert=\bm{U}^{\mathsf{T}}\bm{b}$. Furthermore, since
@@
Now, observe that $B_{i}\leq\lVert\bm{U}^{\mathsf{T}}\bm{S}^{\mathsf{T}}\bm{S}\bm{U}\rVert A_{i}$ and $C_{i}\leq\lVert\bm{I}-\bm{U}^{\mathsf{T}}\bm{S}^{\mathsf{T}}\bm{S}\bm{U}\rVert A_{i}$. Moreover, we have $\lVert\bm{U}^{\mathsf{T}}\bm{S}^{\mathsf{T}}\bm{S}\bm{U}\rVert\leq\eta_{i}$, where $\eta_{i}$ is defined in Eq. 2.3 (and depends on the distribution $\mathcal{P}_{i}$). Therefore,
@@
where we used Cauchy-Schwarz inequality in the last step. Now, note that $\sqrt{\mathbb{E}[A_{i}^{4}]}\geq\mathbb{E}[A_{i}^{2}]$ by Jensen’s inequality, and thus, $\sqrt{\mathbb{E}[A_{i}^{4}]}/\mathbb{E}[A_{i}^{2}]=O(1)$ implies $\sqrt{\mathbb{E}[A_{i}^{4}]}/\mathbb{E}[A_{i}^{2}]=\Theta(1)$. It follows from Eq. A.8 and Eq. A.14 that
@@
Then, denoting $\bm{y}^{*}=\operatornamewithlimits{argmin}_{\bm{y}}\lVert\bm{U}\bm{y}-\bm{b}\rVert$, from Eq. A.23 and Eq. A.4, we obtain
@@
Informally, Eq. D.8 says that (the square-root of) the fourth “central moment" is comparable to the mean squared error of the estimator. This requirement actually holds for the simple statistical task of estimating the mean of a scalar random variable. Indeed, if $X_{1},\dotsc,X_{m}$ are iid copies of a random variable $X$ with $\mathbb{E}[X^{4}]<\infty$, then $\widehat{X}=\sum_{i=1}^{m}X_{i}/m$ is an unbiased estimator of $\mathbb{E}[X]$ satisfying
@@
for all $m$. Motivated by this observation, we expect Eq. D.8 to hold in practice, though this may be difficult to verify rigorously. Also note that while Eq. D.9 gives an expression for ratio of mean squared errors for the regression estimator in the finite-sample regime, the number of samples needs to be large enough so that we can ignore the correction term.
Now, we specialize Theorem D.1 to Shapley value estimation.
@@
Let $\bm{\phi}^{*}$ denote the true Shapley value vector and $\alpha$ as in Theorem 2.1. Given $\lambda\in\mathbb{R}$, define $\bm{b}_{\lambda}$ as in Theorem 2.2. For $i=1,2$, given $m\in\mathbb{N}$ samples from the sampling distribution $\mathcal{P}_{i}$, denote $\bm{\phi}^{\textup{M}}_{\lambda}(\mathcal{P}_{i})$ and $\bm{\phi}^{\textup{R}}_{\lambda}(\mathcal{P}_{i})$ to be the matrix multiplication estimator and regression estimator, respectively. Then, for all $\lambda\in\mathbb{R}$, we have the following results.
@@
Denote $\bm{S}$ to be $m\times 2^{d}-2$ sketching matrix obtained by sampling with replacement according to appropriate sampling probability. Let $\bm{U}$ and $\bm{Q}$ be defined as in Theorem 2.1.
1. From Eq. A.45, we know that $\lVert\bm{\phi}^{\textup{M}}_{\lambda}-\bm{\phi}^{*}\rVert=\lVert\bm{U}^{\mathsf{T}}\bm{S}^{\mathsf{T}}\bm{S}\bm{b}_{\lambda}-\bm{U}^{\mathsf{T}}\bm{b}_{\lambda}\rVert$. Furthermore, from Theorem 2.1, we have $\lVert\bm{U}^{\mathsf{T}}\bm{b}_{\lambda}\rVert=\lVert\bm{Q}\bm{U}^{\mathsf{T}}\bm{b}_{\lambda}\rVert=\lVert\bm{\phi}^{*}-\alpha\bm{1}\rVert$. Then, the result follows from Theorem D.1.
2. From Eq. A.42, we have $\lVert\bm{\phi}^{\textup{R}}_{\lambda}-\bm{\phi}^{*}\rVert=\lVert(\bm{S}\bm{U})^{+}\bm{S}\bm{b}_{\lambda}-\bm{U}^{\mathsf{T}}\bm{b}_{\lambda}\rVert=\min_{\bm{y}}\lVert\bm{S}\bm{U}\bm{y}-\bm{S}\bm{b}_{\lambda}\rVert$. Then, the result follows from Theorem D.1. ∎
The results of this section shows that while the theoretical guarantees derived in Theorem 2.2 only give upper bounds on the sample complexity, the quantity $\gamma$ appearing in this theorem in fact determines the finite-sample/asymptotic behavior of the mean squared errors, as shown in Corollary D.2. Therefore, as long as our metric of performance is the mean squared error, we can directly compare the performance of different sampling schemes by comparing the corresponding values of $\gamma$.
@@ Appendix E Adversarial example
In this section, we develop an adversarial example that help us separate the performance (in terms of the mean squared error) of $\ell_{2}$-squared sampling, kernel weight sampling, and modified $\ell_{2}$ sampling. The main intuition for construction such adversarial examples comes from Corollary A.8 and Corollary D.2, where we compare the value of $\gamma(\bm{z})$ (see Eq. 2.3) for different sampling strategies. The vector $\bm{z}$ is either equal to $\bm{b}_{\lambda}$ or $(\bm{I}-\bm{U}\bm{U}^{T})\bm{b}_{\lambda}$ as in Theorem 2.2. For ease of comparison, in our adversarial example, we will construct a model for which $\bm{b}_{\lambda}=(\bm{I}-\bm{U}\bm{U}^{T})\bm{b}_{\lambda}$, and the lower bounds in Corollary A.8 are saturated up to constant factors. For simplicity, we fix $\lambda=(v([d])-v(\emptyset))/d=\alpha$, as done in previous studies [lundberg2017unified, musco2025provably].
We now construct an example for which we can provably show better theoretical guarantees for modified $\ell_{2}$ sampling and kernel weight sampling compared to $\ell_{2}$-squared sampling. To that end, define the function $f\colon\mathbb{R}^{d}\to\mathbb{R}$, which is our model acting on $d$-dimensional input data, as $f(x)=g(\sum_{i=1}^{d}h(x_{i}))$, where $g$ and $h$ are real-valued functions to be chosen below. While there is a reasonable freedom in defining the functions $h$ and $g$, we choose these judiciously to allows us to theoretically compute the Shapley values. Given a parameter $\epsilon_{0}\in(0,1)$, we define
@@
On the other hand, given parameters $n\in\mathbb{N}$ (independent of $d$) and $\xi,\chi\in\mathbb{R}$, we define $g\colon\mathbb{R}\to\mathbb{R}$ as
@@
For the model $f(\bm{x})=g(\sum_{i=1}^{d}h(x_{i}))$, where $h$ is given in Eq. E.1 and $g$ is given in Eq. E.2, baseline $\bm{y}=\bm{0}$, and explicand $\bm{x}=\bm{1}$, we have $\bm{\phi}^{*}=\chi\bm{1}$ and
@@
First, we compute $\bm{b}_{\lambda}$ and show that $(\bm{I}-\bm{U}\bm{U}^{\mathsf{T}})\bm{b}_{\lambda}=\bm{b}_{\lambda}$. For a given subset $S$ of $[d]$, define $\bm{x}^{S}\in\mathbb{R}^{d}$ as $x^{S}_{i}=x_{i}$ if $i\in S$ and $x^{S}_{i}=y_{i}$ if $y\notin S$. Then, from the definition of $f$, it follows that for all $S\subseteq[d]$, we have $v(S)=f(\bm{x}^{S})=g(|S|)$. By construction, we have $v([d])=\chi d$ and $v(\emptyset)=0$. Since $v(S)$ depends only on the size of the subset $S$, by Eq. 1.1, we have that $\bm{\phi}^{*}=\phi_{0}\bm{1}$ for some constant $\phi_{0}^{*}\in\mathbb{R}$. Then, the constraint $\bm{1}^{T}\bm{\phi}^{*}=v([d])-v(\emptyset)$ gives $\phi_{0}^{*}=(v[d]-v(\emptyset))/d=\chi$. Thus, for this example, we have $\lambda=\alpha=\chi$. Since $\bm{\phi}^{*}=\chi\bm{1}=\alpha\bm{1}$, from Theorem 2.1, we have $\bm{Q}\bm{U}^{\mathsf{T}}\bm{b}_{\lambda}=\bm{0}$, and therefore, $\bm{U}\bm{U}^{\mathsf{T}}\bm{b}_{\lambda}=(\bm{U}\bm{Q}^{\mathsf{T}})(\bm{Q}\bm{U}^{\mathsf{T}})\bm{b}_{\lambda}=\bm{0}$. It follows that $(\bm{I}-\bm{U}\bm{U}^{T})\bm{b}_{\lambda}=\bm{b}_{\lambda}$.
Next, we compute $\left\lVert\bm{b}_{\lambda}\right\rVert^{2}$, $\lVert\bm{H}\bm{b}_{\lambda}\rVert^{2}$, and $\lVert\sqrt{\bm{H}}\bm{b}_{\lambda}\rVert^{2}$ (see Corollary A.9). Since $v$ depends only on the size of the subset and $\lambda=\chi$, we obtain
@@
Therefore, by Corollary A.9, we have
@@
Denote $\mathcal{P}_{\ell_{2}^{2}}$, $\mathcal{P}_{\textnormal{ker}}$, and $\mathcal{P}_{\textnormal{m-}\ell_{2}}$ to be the sampling distributions for $\ell_{2}^{2}$-squared, kernel, and modified $\ell_{2}$ weights, respectively. Then, for the model $f(\bm{x})=g(\sum_{i=1}^{d}h(x_{i}))$, where $h$ is given in Eq. E.1 and $g$ is given in Eq. E.2, baseline $\bm{y}=\bm{0}$, and explicand $\bm{x}=\bm{1}$, we have (using $m$ samples, drawn with replacement)
@@
The expressions for the ratio of mean squared errors for the regression estimator hold under the technical assumption Eq. D.24 stated in Corollary D.2.
@@
This follows by directly substituting the results of Proposition E.1 in Corollary D.2. ∎
This example shows that modified $\ell_{2}$ gives an advantage over leverage scores by a factor of $\sqrt{d}$. On the other hand, kernel weights give a factor of $d/\log(d)$ advantage over leverage scores, while a factor of $\sqrt{d}/\log(d)$ advantage over modified $\ell_{2}$. These saturate the lower bounds in Corollary A.8. Since we have the analytical expressions for $\gamma$ for the adversarial example studied in this section, in Fig. 4, we plot the ratio of $\gamma$ for different the sampling distributions using these expressions.
Figure 4: Ratio of mean-squared errors (MSE) as a function of the dimension for different sampling strategies for the adversarial model in Appendix E (computed analytically from expressions for $\gamma$). The matrix multiplication estimator and regression estimator have (almost) the same MSE ratio for this model (see Corollary E.2). For $\ell_{2}$-squared v/s kernel (solid) and modified $\ell_{2}$ vs kernel (dashed), kernel weights give an advantage by a factor of $\tilde{O}(d)$ and $\tilde{O}(\sqrt{d})$ respectively. On the other hand, for modified $\ell_{2}$ v/s $\ell_{2}$-squared (long dashed), modified $\ell_{2}$ outperforms $\ell_{2}$-squared by a factor of $O(\sqrt{d})$.

### v2

@@ Appendix D Ratio of mean squared errors
Given an $r\times q$ matrix $\bm{U}$ with orthonormal columns, and an $r$-dimensional vector $\bm{b}$, suppose that we want to estimate $\bm{U}^{\mathsf{T}}\bm{b}$ using matrix-vector multiplication estimator (see Theorem˜A.1) and $\operatornamewithlimits{argmin}_{\bm{y}}\left\lVert\bm{U}\bm{y}-\bm{b}\right\rVert^{2}$ using a regression estimator (see Theorem˜A.5). Given a sampling distribution $\mathcal{P}$ over $[r]$ and a fixed number of samples $m$, denote $\bm{X}_{m}^{\textup{M}}(\mathcal{P})$, $\bm{X}_{m}^{\textup{R}}(\mathcal{P})$ to be the matrix-vector multiplication estimator and regression estimator for $\bm{U}^{\mathsf{T}}\bm{b}$, respectively.
Given two sampling distributions $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$, denote $\gamma_{1}$ and $\gamma_{2}$ to be the values of $\gamma$ as defined in Eq.˜2.3 with respect to distributions $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$, respectively. Fix the number of samples $m\in\mathbb{N}$. Then, we have the following results.
@@
Let $\bm{S}$ be an $m\times r$ sketch matrix (for sampling with replacement) as defined in Eq.˜A.2 with respect to the distribution $\mathcal{P}$. Then, the matrix-vector multiplication estimator is $\bm{X}_{m}^{\textup{M}}(\mathcal{P})=\bm{U}^{\mathsf{T}}\bm{S}^{\mathsf{T}}\bm{S}\bm{b}$, while the regression estimator is $\bm{X}_{m}^{\textup{R}}(\mathcal{P})=\operatornamewithlimits{argmin}_{\bm{y}}\lVert\bm{S}\bm{U}\bm{y}-\bm{S}\bm{b}\rVert^{2}$.
1. Matrix-Vector Multiplication: Since $\mathbb{E}[\|\bm{X}_{m}^{\textup{M}}(\mathcal{P})-\bm{U}^{\mathsf{T}}\bm{b}\|^{2}]$ is the variance of $\bm{X}_{m}^{\textup{M}}$ using $m$ samples, from Eq.˜A.4, we have
@@
for $i=1,2$, from which we obtain Eq.˜D.1.
2. Regression: Observe that $\operatornamewithlimits{argmin}_{\bm{y}}\lVert\bm{U}\bm{y}-\bm{b}\rVert^{2}=\bm{U}^{\mathsf{T}}\bm{b}$. Furthermore, since
@@
Now, observe that $B_{i}\leq\lVert\bm{U}^{\mathsf{T}}\bm{S}^{\mathsf{T}}\bm{S}\bm{U}\rVert A_{i}$ and $C_{i}\leq\lVert\bm{I}-\bm{U}^{\mathsf{T}}\bm{S}^{\mathsf{T}}\bm{S}\bm{U}\rVert A_{i}$. Moreover, we have $\lVert\bm{U}^{\mathsf{T}}\bm{S}^{\mathsf{T}}\bm{S}\bm{U}\rVert\leq\eta_{i}$, where $\eta_{i}$ is defined in Eq.˜2.3 (and depends on the distribution $\mathcal{P}_{i}$). Therefore,
@@
where we used Cauchy-Schwarz inequality in the last step. Now, note that $\sqrt{\mathbb{E}[A_{i}^{4}]}\geq\mathbb{E}[A_{i}^{2}]$ by Jensen’s inequality, and thus, $\sqrt{\mathbb{E}[A_{i}^{4}]}/\mathbb{E}[A_{i}^{2}]=O(1)$ implies $\sqrt{\mathbb{E}[A_{i}^{4}]}/\mathbb{E}[A_{i}^{2}]=\Theta(1)$. It follows from Eq.˜A.9 and Eq.˜A.15 that
@@
Then, denoting $\bm{y}^{*}=\operatornamewithlimits{argmin}_{\bm{y}}\lVert\bm{U}\bm{y}-\bm{b}\rVert^{2}$, from Eq.˜A.24 and Eq.˜A.4, we obtain
@@
Informally, Eq.˜D.2 says that (the square-root of) the fourth “central moment" is comparable to the mean squared error of the estimator. This requirement actually holds for the simple statistical task of estimating the mean of a scalar random variable. Indeed, if $X_{1},\dotsc,X_{m}$ are iid copies of a random variable $X$ with $\mathbb{E}[X^{4}]<\infty$, then $\widehat{X}=\sum_{i=1}^{m}X_{i}/m$ is an unbiased estimator of $\mathbb{E}[X]$ satisfying
@@
for all $m$. Motivated by this observation, we expect Eq.˜D.2 to hold in practice, though this may be difficult to verify rigorously. Also note that while Eq.˜D.3 gives an expression for ratio of mean squared errors for the regression estimator in the finite-sample regime, the number of samples needs to be large enough so that we can ignore the correction term.
Now, we specialize Theorem˜D.1 to Shapley value estimation.
@@
Let $\bm{\phi}^{*}$ denote the true Shapley value vector and $\alpha$ as in Theorem˜2.1. Given $\lambda\in\mathbb{R}$, define $\bm{b}_{\lambda}$ as in Theorem˜2.2. For $i=1,2$, given $m\in\mathbb{N}$ samples from the sampling distribution $\mathcal{P}_{i}$, denote $\bm{\phi}^{\textup{M}}_{\lambda}(\mathcal{P}_{i})$ and $\bm{\phi}^{\textup{R}}_{\lambda}(\mathcal{P}_{i})$ to be the matrix-vector multiplication estimator and regression estimator, respectively. Then, for all $\lambda\in\mathbb{R}$, we have the following results.
@@
Denote $\bm{S}$ to be $m\times 2^{d}-2$ sketching matrix obtained by sampling with replacement according to appropriate sampling probability. Let $\bm{U}$ and $\bm{Q}$ be defined as in Theorem˜2.1.
1. From Eq.˜A.47, we know that $\lVert\bm{\phi}^{\textup{M}}_{\lambda}-\bm{\phi}^{*}\rVert=\lVert\bm{U}^{\mathsf{T}}\bm{S}^{\mathsf{T}}\bm{S}\bm{b}_{\lambda}-\bm{U}^{\mathsf{T}}\bm{b}_{\lambda}\rVert$. Furthermore, from Theorem˜2.1, we have $\lVert\bm{U}^{\mathsf{T}}\bm{b}_{\lambda}\rVert=\lVert\bm{Q}\bm{U}^{\mathsf{T}}\bm{b}_{\lambda}\rVert=\lVert\bm{\phi}^{*}-\alpha\bm{1}\rVert$. Then, the result follows from Theorem˜D.1.
2. From Eq.˜A.44, we have $\lVert\bm{\phi}^{\textup{R}}_{\lambda}-\bm{\phi}^{*}\rVert=\lVert(\bm{S}\bm{U})^{+}\bm{S}\bm{b}_{\lambda}-\bm{U}^{\mathsf{T}}\bm{b}_{\lambda}\rVert=\min_{\bm{y}}\lVert\bm{S}\bm{U}\bm{y}-\bm{S}\bm{b}_{\lambda}\rVert$. Then, the result follows from Theorem˜D.1. ∎
The results of this section shows that while the theoretical guarantees derived in Theorem˜2.2 only give upper bounds on the sample complexity, the quantity $\gamma$ appearing in this theorem in fact determines the finite-sample/asymptotic behavior of the mean squared errors, as shown in Corollary˜D.2. Therefore, as long as our metric of performance is the mean squared error, we can directly compare the performance of different sampling schemes by comparing the corresponding values of $\gamma$.
@@ Appendix E Adversarial example
In this section, we develop an adversarial example that help us separate the performance (in terms of the mean squared error) of $\ell_{2}$-squared sampling, kernel weight sampling, and modified $\ell_{2}$ sampling. The main intuition for construction such adversarial examples comes from Corollary˜A.9 and Corollary˜D.2, where we compare the value of $\gamma(\bm{z})$ (see Eq.˜2.3) for different sampling strategies. The vector $\bm{z}$ is either equal to $\bm{b}_{\lambda}$ or $(\bm{I}-\bm{U}\bm{U}^{T})\bm{b}_{\lambda}$ as in Theorem˜2.2. For ease of comparison, in our adversarial example, we will construct a model for which $\bm{b}_{\lambda}=(\bm{I}-\bm{U}\bm{U}^{T})\bm{b}_{\lambda}$, and the lower bounds in Corollary˜A.9 are saturated up to constant factors. For simplicity, we fix $\lambda=(v([d])-v(\emptyset))/d=\alpha$, as done in previous studies [LL17, MW25].
We now construct an example for which we can provably show better theoretical guarantees for modified $\ell_{2}$ sampling and kernel weight sampling compared to $\ell_{2}$-squared sampling. To that end, define the function $f\colon\mathbb{R}^{d}\to\mathbb{R}$, which is our model acting on $d$-dimensional input data, as $f(x)=g(\sum_{i=1}^{d}h(x_{i}))$, where $g$ and $h$ are real-valued functions to be chosen below. While there is a reasonable freedom in defining the functions $h$ and $g$, we choose these judiciously in order to theoretically compute the Shapley values. Given a parameter $\epsilon_{0}\in(0,1)$, we define
@@
Furthermore, given parameters $n\in\mathbb{N}$ (independent of $d$) and $\xi,\chi\in\mathbb{R}$, we define $g\colon\mathbb{R}\to\mathbb{R}$ as
@@
For the model $f(\bm{x})=g(\sum_{i=1}^{d}h(x_{i}))$, where $h$ is given in Eq.˜E.1 and $g$ is given in Eq.˜E.2, baseline $\bm{y}=\bm{0}$, and explicand $\bm{x}=\bm{1}$, we have $\bm{\phi}^{*}=\chi\bm{1}$ and
@@
First, we compute $\bm{b}_{\lambda}$ and show that $(\bm{I}-\bm{U}\bm{U}^{\mathsf{T}})\bm{b}_{\lambda}=\bm{b}_{\lambda}$. For a given subset $S$ of $[d]$, define $\bm{x}^{S}\in\mathbb{R}^{d}$ as $x^{S}_{i}=x_{i}$ if $i\in S$ and $x^{S}_{i}=y_{i}$ if $y\notin S$. Then, from the definition of $f$, it follows that for all $S\subseteq[d]$, we have $v(S)=f(\bm{x}^{S})=g(|S|)$. By construction, we have $v([d])=\chi d$ and $v(\emptyset)=0$. Since $v(S)$ depends only on the size of the subset $S$, by Eq.˜1.1, we have that $\bm{\phi}^{*}=\phi_{0}\bm{1}$ for some constant $\phi_{0}^{*}\in\mathbb{R}$. Then, the constraint $\bm{1}^{T}\bm{\phi}^{*}=v([d])-v(\emptyset)$ gives $\phi_{0}^{*}=(v[d]-v(\emptyset))/d=\chi$. Thus, for this example, we have $\lambda=\alpha=\chi$. Since $\bm{\phi}^{*}=\chi\bm{1}=\alpha\bm{1}$, from Theorem˜2.1, we have $\bm{Q}\bm{U}^{\mathsf{T}}\bm{b}_{\lambda}=\bm{0}$, and therefore, $\bm{U}\bm{U}^{\mathsf{T}}\bm{b}_{\lambda}=(\bm{U}\bm{Q}^{\mathsf{T}})(\bm{Q}\bm{U}^{\mathsf{T}})\bm{b}_{\lambda}=\bm{0}$. It follows that $(\bm{I}-\bm{U}\bm{U}^{T})\bm{b}_{\lambda}=\bm{b}_{\lambda}$.
Next, we compute $\left\lVert\bm{b}_{\lambda}\right\rVert^{2}$, $\lVert\bm{H}\bm{b}_{\lambda}\rVert^{2}$, and $\lVert\sqrt{\bm{H}}\bm{b}_{\lambda}\rVert^{2}$ (see Corollary˜A.10). Since $v$ depends only on the size of the subset and $\lambda=\chi$, we obtain
@@
Therefore, by Corollary˜A.10, we have
@@
Denote $\mathcal{P}_{\ell_{2}^{2}}$, $\mathcal{P}_{\textnormal{ker}}$, and $\mathcal{P}_{\textnormal{m-}\ell_{2}}$ to be the sampling distributions for $\ell_{2}^{2}$-squared, kernel, and modified $\ell_{2}$ weights, respectively. Then, for the model $f(\bm{x})=g(\sum_{i=1}^{d}h(x_{i}))$, where $h$ is given in Eq.˜E.1 and $g$ is given in Eq.˜E.2, baseline $\bm{y}=\bm{0}$, and explicand $\bm{x}=\bm{1}$, we have (using $m$ samples, drawn with replacement)
@@
The expressions for the ratio of mean squared errors for the regression estimator hold under the technical assumption Eq.˜D.18 stated in Corollary˜D.2.
@@
This follows by directly substituting the results of Proposition˜E.1 in Corollary˜D.2. ∎
This example shows that modified $\ell_{2}$ gives an advantage over leverage scores by a factor of $\sqrt{d}$. On the other hand, kernel weights give a factor of $d/\log(d)$ advantage over leverage scores, while a factor of $\sqrt{d}/\log(d)$ advantage over modified $\ell_{2}$. These saturate the lower bounds in Corollary˜A.9. Since we have the analytical expressions for $\gamma$ for the adversarial example studied in this section, in Fig.˜4, we plot the ratio of $\gamma$ for different the sampling distributions using these expressions.
Figure 4: Ratio of mean-squared errors (MSE) as a function of the dimension for different sampling strategies for the adversarial model in Appendix˜E (computed analytically from expressions for $\gamma$). The matrix-vector multiplication estimator and regression estimator have (almost) the same MSE ratio for this model (see Corollary˜E.2). For $\ell_{2}$-squared v/s kernel (solid) and modified $\ell_{2}$ vs kernel (dashed), kernel weights give an advantage by a factor of $\tilde{O}(d)$ and $\tilde{O}(\sqrt{d})$ respectively. On the other hand, for modified $\ell_{2}$ v/s $\ell_{2}$-squared (long dashed), modified $\ell_{2}$ outperforms $\ell_{2}$-squared by a factor of $O(\sqrt{d})$.